A rectangular room is twice as long as it is wide, and its perimeter is 24 meters. Find the length of the room.
Answer by TimothyLamb(4379)

L = 2w
2L + 2w = 24
---
2L + 2w = 24
2L + L = 24
3L = 24
L = 8
---
answer:
length is 8 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php